CH13 Q9

A researcher conducts a two-way ANOVA test with interaction and provides the following ANOVA table a. Find the missing values in the ANOVA table. (Round intermediate calculations to at least 4 decimal places. Round MS to 4 decimal places and F to 3 decimal places.) ANOVA Source of Variation Sample Columns df MS 770.76 12,351.18 55.10 693.74 13,870.78 F Factor B- F Factor A F Interaction- 2 MSB 1MSA- Within 12 MSE- 17 Total
Answer #1

Solution :

From give ANOVA table,

MSB = SSB / df = 770.76 / 2 = 385.38

MSA = SSA / df = 12351.18 / 1 = 12351.18

MSAB = SSAB / df = 55.10 / 2 = 27.55

MSE = 693.74 / 12 = 57.8117

F factor B = MSB / MSE = 385.38 / 57.81 = 6.6661

F factor A =  MSA / MSE = 12351.18 / 57.8117 = 213.6445

interaction = MSAB / MSE = 27.55 / 57.8117 = 0.4765